6. Suppose that in a population of 400,000 people, there are 100,000 people who are either institutionalized or are too young to work. Additionally, there are 60,000 people who choose not to work because they are stay-at-home parents, college students, or retired. In this scenario, if 228,000 people currently are employed full-time, what is the unemployment rate?

a- 2.28%

b- 57%

c- 10%

d- 5%

7. Suppose that in a population of 400,000 people, there are 100,000 people who are either institutionalized or are too young to work. Additionally, there are 60,000 people who choose not to work because they are stay-at-home parents, college students, or retired. In this scenario, what is the size of the labor force?

a- 300,000

b- 400,000

c- 240,000

d- 360,000

8. All of the following are causes of shocks that can cause business cycles, except:

a- Political Events

b- Sudden or Dramatic Changes in Productivity

c- Consistent, Steady Innvation

d- Financial Instability

9. Which is the correct sequencing of the business cycle?

a- Expansion, Recession, Trough, Peak

b- Peak, Expansion, Recession, Trough

c- Expansion, Peak, Recession, Trough

d- Trough, Recession, Expansion, Peak

10. Economists define a recession as:

a- Two or more consecutive quarters or declining real GDP

b- Any reduction in employment that occurs from one period to the next

c- Whenever GDP from one period is lower than the period before it

d- Six months of declining GDP
Answer #1

6. Option D. 5%

Explanation: Unemployment rate = unemployed/labor force * 100 = unemployed/(employed+unemployed) * 100.

Here, unemployed = Population - (Institutionalized people + People choosing not to work + employed) = 400,000 - (100,000 + 60,000 + 228,000) = 400,000 - 388,000 = 12,000.

So, Unemployment rate = unemployed/labor force * 100 = unemployed/(employed+unemployed) * 100 = 12,000/(228,000 + 12,000) *100 = 12,000/240,000 *100 = 12/240 * 100 = 1200/240 = 5%

7. Option C

Explanation: Labor force = employed + unemployed = 228,000 + 12,000 = 240,000.

8. Option C. Consistent, Steady Innovation

Explanation: Consistent and steady innovation does not result in output in a significant way.

9. Option C. Expansion, Peak, Recession, Trough

Explanation: Peak is the highest point of expansion and trough is the lowest point of recession.

10. Option A. Two or more consecutive quarters or declining real GDP

Explanation: Recession takes place when the real GDP declines for 2 consecutive periods.